OpenStudy (anonymous):

Please help! If the diagonal of a parallelogram measures 40 cm, the angle opposite that diagonal is 50°11', and one side is 24 cm, find the other side of the parallelogram.

jimthompson5910 (jim_thompson5910):

A drawing may help 

jimthompson5910 (jim_thompson5910):

Now you can use the law of cosines to solve for x c^2 = a^2 + b^2 - 2ab*cos(C) 40^2 = 24^2 + x^2 - 2*24*x*cos(50 degrees 11 minutes) Keep going to solve for x

OpenStudy (anonymous):

I tried that but you get a quadratic with messy numbers, how do you solve that?

jimthompson5910 (jim_thompson5910):

what did you get

jimthompson5910 (jim_thompson5910):

you would get a quadratic because of the x^2 and the x terms, so you can't avoid that

OpenStudy (anonymous):

0 = x^2 -48*cos(50°11')*x -1024

jimthompson5910 (jim_thompson5910):

ok good, now you can evaluate the cosine of 50°11' to get some number keep in mind 50°11' = 50 + 11/60 50°11' = 50 + 0.18333333333333 50°11' = 50.18333333333333 degrees

OpenStudy (anonymous):

So I guess I can simplify it to 0 = x^2 - 30.73599154 - 1024. But see, that's kind of messy (and I definitely can't factor that easily!)

jimthompson5910 (jim_thompson5910):

If you can't factor, then use the quadratic formula

OpenStudy (anonymous):

Hmm okay. x^2 - 30.73599154 - 1024 = 0 a = 1, b = -30.73599154 (for now I'll leave it as b), c = -1024  That's not the answer :(